# The Open Proof Corpus: A Large-Scale Study of LLM-Generated Mathematical Proofs

### v1

@@ 6 Limitations
While the OPC represents a significant advancement in the development and evaluation of LLM proof-generation capabilities, it is not without limitations. First, since most problems in the OPC are sourced from publicly available mathematical competitions, there is a risk of data contamination. Some of these problems may have been encountered during the training of the models. However, since the main purpose of the OPC is not related to benchmarking but rather to provide a large, human-validated dataset of LLM-generated proofs, this issue is less critical. Second, the majority of problems in the OPC are derived from high school level competitions. As a result, the dataset does not cover more advanced mathematical domains, such as undergraduate or research-level mathematics.
7 Future Work
The OPC offers a valuable resource for advancing research in proof generation with LLMs, supporting both the training and evaluation of models across a diverse set of mathematical problems. Our analysis also highlights several critical gaps and challenges that warrant further research. First, the significant disparity between formal and natural language proof generation—outlined in Section˜5.3—points to the need for more effective strategies to bridge this divide. Second, the OPC demonstrates that current benchmarks, which often rely solely on final answers, fail to capture the full complexity and quality of generated proofs, as discussed in Section˜5.4. This underscores the need for the development of a scalable benchmarking pipeline tailored to proof generation tasks. Finally, while our results show that best-of-n sampling strategies can meaningfully improve proof quality (see Section˜5.5), further research is required to better understand and optimize these methods.
@@ 8 Conclusion
In this work, we introduced the Open Proof Corpus (OPC), a large-scale, human-validated dataset of over 5,0005{,}000 LLM-generated proofs across more than 1,0001{,}000 problems. Designed to support training, evaluation, and benchmarking, the OPC provides a robust foundation for advancing research in automated proof generation. Using the OPC, we conducted a thorough evaluation of current LLM capabilities in mathematical reasoning. Our analysis reveals several key findings: (1) modern LLMs can generate a significant number of correct proofs; (2) some models approach human-level performance in judging proof correctness; (3) natural language proof generation substantially outperforms formal proof generation; (4) final-answer benchmarks are a poor proxy for proof quality; and (5) best-of-n sampling strategies significantly improve proof-generation accuracy. Using the OPC, we also trained an open-source model that achieves $88.1\%$ accuracy in judging proof correctness, matching the performance of the best closed-source model, Gemini-2.5-Pro.
@@ Appendix A Competitions in the OPC
The OPC contains over 1000 problems that were sourced from national and international competitions of varying difficulty. In Table˜4, we present the problem and sample distribution for each. We also include the following additional information:
@@
  • Source: we link the source, from which we obtained the problems. Any source that is not publicly available was marked as "Private".
@@ Appendix B Swiss Ranking Methodology
Appendix C Qualitative Observations
@@
Many models sometimes make mistakes when manipulating inequalities. A frequent issue is the improper reversal of inequality signs during intermediate steps without mathematical justification. This results in incorrect final answers despite seemingly coherent reasoning. Below is an example of such a mistake made by DeepSeek-R1, where the incorrect sign reversal is marked with "**".
[Uncaptioned image]  DeepSeek-R1 – Baltic Way 2004 P3
@@
Appendix D Grading Interface and Instructions
This appendix outlines the grading interface and the accompanying instructions provided to judges. The full interface and documentation can be accessed online at https://judgeproofs.xyz/sample and https://judgeproofs.xyz/instructions.
@@ Appendix D Grading Interface and Instructions
Each judge received a unique identifier, which served as their login credential on https://judgeproofs.xyz. This ID was used to track grading progress while maintaining judge anonymity in the resulting dataset. To facilitate discussion and resolve ambiguities, a shared communication channel was created between all judges.
@@
The grading interface was designed for clarity and ease of use. Figs.˜8, 9 and 10 illustrate its main components. The left panel contains a navigation bar for switching between problems and competitions. The right panel displays the problem statement and the ground-truth solution, along with options for flagging issues in either. Below, the generated solution is shown, accompanied by an automated summary and potential issues identified by an LLM judge. Judges can then evaluate the solution using a grading form that allows them to:
@@
Appendix E Prompts
In this section, we provide the prompts used for various tasks in the OPC. The prompts are designed to be clear and concise, guiding the LLMs through the proof generation process while ensuring that they understand the requirements for correctness and clarity. In Section˜E.1, we present the prompts used for generating proofs. In Section˜E.2, we provide the prompt used to generate the LLM summary to aid human graders in identifying potential issues in the proof. In Section˜E.3, we present the prompt used for LLMs to judge the correctness of a proof, used in Section˜5.2. In Sections˜E.4, E.5 and E.6, we provide all prompts used for the LLMs in best-of-n sampling, as described in Section˜5.5.
E.1 Proof Generation Prompt
@@
E.2 Issues Interface Prompt
@@ E.3 LLM as Judge Prompt
E.4 LLM as Discrete Judge Prompt
@@
E.5 LLM as Continuous Judge Prompt
@@
E.6 LLM as Rank Judge Prompt

### v2

@@ 6 Limitations
While the OPC represents a significant advancement in the development and evaluation of LLM proof-generation capabilities, it is not without limitations. First, since dataset construction took place before the release of Grok-4 and GPT-5, these models are only included as judges. However, recent benchmarks suggest that these models perform similarly, perhaps slightly better, compared to Gemini-2.5-Pro (Balunović et al., 2025). Therefore, it does not affect the validity of our conclusions. Second, the majority of problems in the OPC are derived from high-school competitions. Thus, the dataset does not cover more advanced mathematical domains, such as research-level mathematics. We further outline potential directions for expanding the OPC in App.˜F.
@@ 7 Conclusion
In this work, we introduced the Open Proof Corpus (OPC), a human-validated dataset comprising over 5,0005{,}000 LLM-generated proofs. Using the OPC, we addressed several open questions in the field, including (1) the gap between natural language and formal proof generation, (2) the relationship between final-answer accuracy and proof correctness, and (3) the effectiveness of best-of-n selection. We also trained a model that achieves $88.1\%$ accuracy in judging proof correctness, matching Gemini-2.5-Pro and approaching the performance of the top model, GPT-5.
Reproducibility Statement
We have included our dataset in the supplementary material, along with detailed descriptions of our methodology and experimental setup to ensure full reproducibility. In addition, we provide our code with step-by-step instructions in the supplementary material, enabling others to replicate our results and train the model from scratch. While the model itself is open-sourced, it is not included in the supplementary materials due to file size constraints. A small portion of the problems included in the dataset were obtained through private correspondence with the competition organizers (SMT 2025, $\approx 2\%$ of the data). They asked that we do not share these problems until they have released them publicly. We will therefore share these problems once they are made public by the competition organizers. Our results remain virtually unchanged if we exclude these problems from our dataset.
@@ Appendix A Competitions in the OPC
The OPC contains over 1000 problems that were sourced from national and international competitions of varying difficulty. In Table˜5, we present the problem and sample distribution for each. We also include the following additional information:
@@
  • Source: we link the source from which we obtained the problems. Any source that is not publicly available was marked as "Private".
@@ Appendix B Swiss Ranking Methodology
Appendix C Additional Results
In this section, we present further results, specifically addressing critical aspects of our evaluation: out-of-distribution performance of our trained model and the potential impact of data contamination.
C.1 OPC-R1-8B performance on OOD problems
During the training of OPC-R1-8B, our dataset primarily consisted of samples from high-school level competitions, aligning with the distribution of the test set. Crucially, undergraduate-level problems were explicitly excluded from training. To evaluate OPC-R1-8B’s generalization capabilities on out-of-distribution data, we evaluated its judging performance on 560 undergraduate-level solution samples sourced from the Putnam competition.
Table 6: LLMs as proof graders on undergraduate-level problems. Cost for running the model on the entire subset is given in USD. Confidence intervals are 95% and computed using the large sample normal approximation.
Judge pass@1 maj@5 Cost
o4-mini $83.8\pm 3.1$ $85.1\pm 2.9$ 48.31
GPT-5 $83.6\pm 3.1$ $84.0\pm 3.0$ 89.42
R1 $78.3\pm 3.4$ $79.7\pm 3.3$ 51.86
OPC-R1-8B $5.0\pm 3.6 77$$7.0\pm 3.5$ N/A
R1-Qwen3-8B $72.1\pm 3.7$ $72.4\pm 3.7$ N/A
Gemini-2.5-Pro $70.9\pm 3.8$ $76.7\pm 3.5$ 160.95
As shown in Table˜6, OPC-R1-8B retains a notable improvement over its base model, despite undergraduate-level examples not being present in its testing set. While frontier models like o4-mini achieve higher performance, OPC-R1-8B is nevertheless competitive, performing on par with Gemini-2.5-Pro. This shows that the OPC data can be applied to mathematical domains beyond its scope.
C.2 Problem Category and Difficulty Affect Incorrect Proof Rates
We analyze the prevalence of incorrect proofs in the MathArena subset of the OPC. In particular, we investigate whether difficulty and problem type correlates with the likelihood of a model generating an incorrect proof. As shown in Fig.˜7, we find that both these factors significantly influence performance. In particular, higher difficulty levels correspond to a greater proportion of incorrect proofs, as illustrated in Fig.˜7(a). Additionally, among the four main problem categories, Geometry problems exhibit the highest rate of incorrect proofs.
((a)) Problems are binned by difficulty level, with higher levels indicating more challenging problems.
((b)) Problems are binned in the four main categories: Algebra, Combinatorics, Geometry, and Number Theory.
Figure 7: Influence of difficulty level and problem category on the prevalence of incorrect proofs in MathArena.
C.3 Data Contamination for Mathematical Proofs
Test-set contamination is a prevalent challenge in evaluating the true reasoning capabilities of models, particularly in mathematics [Balunović et al., 2025, Zhang et al., 2024]. Given that the vast majority of OPC problems are publicly sourced, we conduct one additional experiment to address potential contamination-related concerns.
Competition-based breakdown
First, we outline 2 categories of competitions: Standard and Non-standard. The former includes important competitions in the mathematics community, namely the IMO Shortlist and USAMO. Problems from these widely recognized competitions are more likely to appear in training sets, as they are widely reported on the internet and therefore represent the greatest potential risk for contamination.
Figure 8: Average proof correctness on a sample of the OPC, split into Standard and Non-standard competitions.
On the other hand, the Non-standard category consists of problems from the International Zhautykov Olympiad, the Swiss Mathematical Olympiad, and the Bulgarian seasonal competitions. A significant portion of these problems required translation from their original non-English statements, and they originate from less prevalent sources compared to the IMO and USAMO. Consequently, we anticipate a substantially lower prevalence of contamination within this problem set.
As illustrated in Fig.˜8, model accuracy on Standard problems is notably lower than on their Non-standard counterparts. This observation indicates that problem difficulty is a considerably more dominant factor influencing model performance than the potential presence of test-set contamination. While this experiment alone cannot definitively discount the presence of test-set contamination within the OPC dataset for our Section˜5.1 results, it demonstrates that the models’ problem-solving capabilities extend significantly beyond memorization. This suggests that any contamination effect is likely small and not the primary driver of performance differences.
C.4 Significance Analysis
In Table˜7, we present the performance of various LLMs as proof judges, along with 95% confidence intervals using the large sample normal approximation. In Table˜8, we provide a detailed breakdown of judging accuracy by prover, again with 95% confidence intervals. These tables complement the main results presented in Section˜5.
Table 7: LLMs as proof graders. Cost for running the model on the entire subset is given in USD. Confidence intervals are computed using the large sample normal approximation.
Judge pass@1 maj@5 Cost
Human $90.4\pm 3.4$ - N/A
GPT-5 $89.3\pm 3.5$ $90.8\pm 3.3$ $117.77$
Grok-4 $88.3\pm 3.7$ $89.8\pm 3.5$ $104.42$
Gemini-2.5-Pro $5.4\pm 4.0 88$$8.1\pm 3.7$ $135.47$
OPC-R1-8B $3.8\pm 4.2 88$$8.1\pm 3.7$ N/A
o4-mini $83.8\pm 4.2$ $85.3\pm 4.1$ $29.57$
o3 $83.1\pm 4.3$ $84.3\pm 4.2$ $93.3$
Gemini-2.5-Flash $82.7\pm 4.3$ $86.0\pm 4.0$ $86.95$
Qwen3-235B-A22B $81.8\pm 4.4$ $84.6\pm 4.1$ $3.79$
R1 $80.9\pm 4.5$ $82.6\pm 4.3$ $27.70$
R1-Qwen3-8B $70.7\pm 5.2$ $71.3\pm 5.2$ N/A
Claude-4-Sonnet $70.6\pm 5.2$ $75.0\pm 5.0$ $28.21$
Qwen3-8B $64.4\pm 5.5$ $63.6\pm 5.5$ N/A
GPT-4.1 $61.4\pm 5.6$ $60.8\pm 5.6$ $20.33$
Table 8: Judgement accuracy breakdown by prover with 95% confidence intervals.
Prover Judge Gemini o4 o3 Qwen
Gemini $79.4\pm 4.9$ $86.9\pm 4.1$ $85.9\pm 4.2$ $80.0\pm 4.9$
o4 $87.1\pm 4.1$ $81.3\pm 4.8$ $84.8\pm 4.4$ $81.9\pm 4.7$
o3 $91.6\pm 3.4$ $83.1\pm 4.6$ $76.9\pm 5.2$ $79.1\pm 5.0$
Qwen $80.6\pm 4.9$ $84.1\pm 4.6$ $87.8\pm 4.1$ $84.4\pm 4.5$
Appendix D Experimental Details
In this section, we provide additional details regarding our experimental setup, including model training, hyperparameters, and evaluation protocols.
D.1 Model Training
Training data
As outlined in Section˜5, we split the generic split of the OPC dataset into a train and test set. The training set consists of 1,733 proof samples, while the test set contains 293 proof samples. Importantly, we ensure that no problem statements overlap between the training and test sets, maintaining the integrity of our evaluation.
Reinforcement learning
We fine-tune R1-Qwen3-8B using GRPO [Shao et al., 2024] on the training set using the popular VERL framework [Sheng et al., 2024] with a learning rate of $10^{-6}$, a maximum response length of $14000$ tokens, $10$ rollouts per problem, and a batch size of $16$. We use the same prompt for training as for evaluation, as shown in Section˜I.3. More hyperparameters and training details can be found in the provided code. Importantly, we did not use the test set at any point during training or hyperparameter tuning, and only evaluated the final trained model once on the test set.
D.2 Model Evaluation
Hyperparameters and prompt
For all models, we use the recommended hyperparameters outlined by the model providers. For open models, this usually involves using nucleus sampling with $p=0.95$ and a temperature of $0.6$. For closed models, we use the default settings provided by the API. Model-specific hyperparameters can be found in the provided code. Prompts used for evaluation are shown in App.˜I.
Formal model performance
To accurately compute the performance of Goedel-Prover-V2 [Lin et al., 2025b] on our subset of problems from Tsoukalas et al. [2024], we contacted the authors of the models to obtain the raw results for each problem. This allowed us to compute the exact accuracy on our selected problems, which was not reported in the original paper.
Appendix E Qualitative Observations
@@
Many models sometimes make mistakes when manipulating inequalities. A frequent issue is the improper reversal of inequality signs during intermediate steps without mathematical justification. This results in incorrect final answers despite seemingly coherent reasoning. Below is an example of such a mistake made by R1, where the incorrect sign reversal is marked with "**".
[Uncaptioned image]  R1 – Baltic Way 2004 P3
@@
Appendix F Future Work
The OPC offers a valuable resource for advancing research in proof generation with LLMs, supporting both the training and evaluation of models for this task. Our analysis also highlights several critical gaps and challenges that require further research. First, the significant disparity between formal and natural language proof generation shows the need for more effective strategies to bridge this gap. Second, the OPC shows that current benchmarks fail to capture the full complexity and quality of generated proofs as they often rely solely on final answers. Therefore, the development of a scalable benchmarking pipeline tailored to proof generation tasks is necessary. Finally, while our results show that best-of-n sampling strategies can meaningfully improve proof quality, further research is required to better understand and optimize these methods.
Appendix G Statement for the Use of Large Language Models
Beyond being the subject of our research, LLMs were used only as assistants to improve the clarity and quality of writing. They were not involved in aiding our research methodology, ideation, or for discovering related work.
Appendix H Grading Interface and Instructions
This appendix outlines the grading interface and the accompanying instructions provided to judges. The full interface and documentation can be accessed and reviewed in our supplementary material.
@@ Appendix H Grading Interface and Instructions
Each judge received a unique identifier, which served as their login credential on our website. This ID was used to track grading progress while maintaining judge anonymity in the resulting dataset. To facilitate discussion and resolve ambiguities, a shared communication channel was created between all judges.
@@
The grading interface was designed for clarity and ease of use. Figs.˜9, 10 and 11 illustrate its main components. The left panel contains a navigation bar for switching between problems and competitions. The right panel displays the problem statement and the ground-truth solution, along with options for flagging issues in either. Below, the generated solution is shown, accompanied by an automated summary and potential issues identified by an LLM judge. Judges can then evaluate the solution using a grading form that allows them to:
@@
Appendix I Prompts
In this section, we provide the prompts used for various tasks in the OPC. The prompts are designed to be clear and concise, guiding the LLMs through the proof generation process while ensuring that they understand the requirements for correctness and clarity. In Section˜I.1, we present the prompts used for generating proofs. In Section˜I.2, we provide the prompt used to generate the LLM summary to aid human graders in identifying potential issues in the proof. In Section˜I.3, we present the prompt used for LLMs to judge the correctness of a proof, used in Section˜5.2. In Sections˜I.5, I.6 and I.7, we provide all prompts used for the LLMs in best-of-n sampling, as described in Section˜5.5.
I.1 Proof Generation Prompt
@@
I.2 Issues Interface Prompt
@@ I.3 LLM as Judge Prompt
I.4 LLM as Judge Prompt with Ground Truth Solution
@@
I.5 LLM as Discrete Judge Prompt
@@
I.6 LLM as Continuous Judge Prompt